# DiffSeg30k: A Multi-Turn Diffusion Editing Benchmark for Localized AIGC Detection

### v1

@@ 1 Introduction
In this paper, we introduce DiffSeg30k, a publicly available dataset comprising 30k diffusion-edited images with pixel-level annotations, specifically aimed at facilitating the detection and localization of diffusion-based edits. To reflect realistic, practical scenarios, DiffSeg30k incorporates four key features: 1) In-the-wild images—we collect natural images from COCO dataset [coco] to ensure real-world content diversity. In addition, we leverage COCO-derived prompts with diffusion models to generate complementary AI-based images; 2) Diverse diffusion models—local editing is conducted using eight state-of-the-art diffusion models; 3) Realistic user-driven edits—an automated annotation pipeline leveraging vision-language models (VLMs) [Qwen2.5-VL] identifies semantically meaningful regions and generates context-aware editing prompts, including object addition, removal, and attribute modification; and 4) Multi-turn editing—each image undergoes up to three sequential edits using different diffusion models, simulating realistic scenarios in which online-shared images are iteratively modified (see Figure LABEL:fig:teaser). Beyond pixel-level annotations of edited areas, the specific diffusion model used for each edit is annotated, defining a novel semantic segmentation task for fine-grained localization that simultaneously identifies the edited regions and the corresponding editing models. Table 1 compares DiffSeg30k with existing benchmarks containing local editing data.
@@ 2.2 AIGC Detection
The task of AIGC detection involves identifying content created by generative AI models such as GANs and diffusion models. Current datasets [zhu2023genimage, defake, bird2024cifake, verdoliva20222022, deepart, he2021forgerynet, guo2023hierarchical, miao2025ddl] and methods [cozzolino2024raising, cnnspot, zhong2023patchcraft, cozzolino2024zero, wang2025scaling] predominantly focus on classifying entire images. Existing datasets can be categorized into two types according to their image content: facial images [he2021forgerynet, yang2019exposing, wang2019fakespotter, dang2020detection, gandhi2020adversarial, miao2025ddl] and general images [zhu2023genimage, defake, bird2024cifake, verdoliva20222022, deepart]. Several datasets [guillaro2023trufor, wfake, epstein2023online, liu2024hfnet] include limited numbers ($\leq 3$) of diffusion models and only involving single local edit per image, as detailed in Table 1. In contrast, our proposed DiffSeg30k dataset focuses on general images, covers eight different diffusion models, and incorporates multi-turn editing by different models on each image. This comprehensive dataset provides systematic support for more general semantic segmentation tasks beyond conventional first-classify-then-binary-segmentation paradigm [guo2023hierarchical] used in single-turn editing datasets.
@@ 3.1 Dataset construction
We developed an automated image editing and annotation pipeline based on vision-language models (VLM), illustrated in Figure 2. The pipeline comprises two stages: 1) generating candidate editing regions and 2) creating suitable editing prompts for these regions, followed by image editing.
@@
Stage 1—generating candidate editing regions. As shown in Figure 2, given an arbitrary image, we first use Qwen2.5-VL [Qwen2.5-VL] to identify distinct objects present in the image. The identified object categories are then provided to Grounded-SAM [groundedsam] to generate corresponding masks for each object. Subsequently, we calculate the Intersection-over-Union (IoU) among object masks and eliminate redundant masks with IoU greater than 70% to avoid repetitive editing of the same object. Finally, we randomly select between one to three object masks for editing.
@@
Stage 2—editing prompt generation & inpainting. As depicted in Figure 2, each selected object mask from Stage 1 is paired with the original image and fed into the Qwen2.5-VL model, prompting it to generate contextually appropriate editing prompts. For inpainting, the generated editing prompt, corresponding object mask, and the original image are simultaneously provided to a diffusion model to perform localized inpainting. For multi-turn editing, the edited image can form a new pair with the remaining object masks, repeating the prompt generation and inpainting processes. Thus, we obtain images edited sequentially up to three times, depending on the number of masks selected in Stage 1.
@@
Data diversity & balance. To ensure the diversity and balance of the edited images, we specifically considered the following six aspects: 1) Balanced Candidate Object Types: We prompted the VLM to prioritize selecting humans, as human-centric edits are significant in practical applications, yet VLMs tend to under-select humans naturally. 2) Balanced Candidate Mask Area: We encouraged VLM to prioritize larger objects for editing, counteracting the empirical tendency of VLM to select smaller objects. 3) Balanced Edit Types: During editing prompt generation, we prompted VLM to randomly choose from three edit types: attribute changes, object additions, and object removals. For object additions, we followed the approach of SEED-Data-Edit [ge2024seed], first removing an object and then adding a new object in the same position to ensure natural placement. 4) Balanced Edit Models: For inpainting, we randomly selected one diffusion model from eight state-of-the-art diffusion models to execute an editing prompt. 5) Enhanced Multi-turn Editing: We manually increased the frequency of multi-turn editing, with the first, second, and third turns occurring at an approximate 1:4:5 ratio, as multi-turn editing subsumes single-turn cases and presents a more challenging scenario. 6) Balanced Real and AI Bases: Approximately half of the base images were real, collected from COCO, while the other half were AI-generated using COCO prompts with random diffusion models.
@@ 3.3 Dataset statistics
We conducted a statistical analysis of DiffSeg30k, illustrated in Figure 3. It reveals that smaller objects are edited more frequently, whereas larger objects tend to be edited less often, with editing frequency smoothly decreasing as object area increases. Objects such as food, vehicles, furniture, nature scenery, animals, humans, and signs are among the most frequently edited. The dataset maintains a balanced distribution across editing models and base image distribution. Figure 4 shows some editing results.
@@ 4.1 Experimental setup
Training details. Due to the lack of existing baseline segmentation models on general AIGC localization to handle semantic segmentation, we trained two CNN baselines based on classic semantic segmentation frameworks: FCN-8s [long2015fully] and Deeplabv3+[chen2017rethinking, chen2018encoder] with ResNet50[he2016deep] and 1 transfomer baseline SegFormer-B2 [xie2021segformer]. We evaluated these models on two tasks: binary segmentation (localizing edited areas) and semantic segmentation (localizing edited areas while distinguishing among different editing models). The dataset is split into training and validation sets in an 8:2 ratio, and all results are reported on the validation set. Models are trained on 512×512 resolution (for acceleration) using default hyperparameters from open-source implementations. We additionally include 5k unedited real COCO images in the training set to help suppress false positives on normal images. Training FCN-8s, SegFormer-B2, and Deeplabv3+ takes approximately 8, 6, and 3 hours, respectively, on an NVIDIA 4090 GPU.
@@ 4.2 Results
Model capacity matters. As shown in Table 2, the older FCN-8s struggles with both binary and semantic localization of diffusion-edited regions. In contrast, more powerful segmentation models Deeplabv3+ and SegFormer , performs substantially better, achieving strong binary localization with an mIoU of 0.974 and 0.961. These results highlight the importance of model representational capacity.
Semantic segmentation is challenging. Simultaneously localizing edited regions and attributing them to the correct diffusion model is substantially more difficult than binary localization. Even for the best performed model SegFormer, the mIoU for semantic segmentation drops to 0.825, leaving significant room for improvement. For Deeplabv3+ and FCN, this number further drops significantly to 0.76 and 0.23. Figure 5 shows the confusion matrices: FCN-8s tends to misclassify edits from all models as background pixels, while SegFormer struggles particularly with Flux. Deeplab v3+ struggles with Kolors and SDXL. It also often misclassifies SD2 and Flux. Figure 6 further illustrates representative results. FCN-8s produces largely random outputs in both tasks, whereas SegFormer and Deeplab v3+ achieves more coherent segmentations but occasionally misclassifies the editing model. In the following detailed ablation, we only evaluate Deeplabv3+ and SegFormer.
Baseline models are sensitive to image transformations. We evaluated the robustness of SegFormer and DeepLabv3+ under common image transformations, including JPEG compression and resizing. As shown in Table 3, image transformations substantially degraded the performance of both Deeplabv3+ and SegFormer, with severe mIoU drops—semantic segmentation almost failed under JPEG compression. This highlights the need for improved network architectures or data augmentation strategies to achieve deployable localization models. Interestingly, DeepLabv3+ exhibits partial resilience in binary segmentation settings, while SegFormer demonstrates certain robustness when resized to higher resolutions.
Segmentation models achieve strong cross-generator generalization. Table 4 presents the cross-generator generalization results of SegFormer and Deeplabv3+. Training was conducted on images edited by six diffusion models and tested on edits from two unseen models, with three different combinations evaluated. Both SegFormer and Deeplabv3+ demonstrate strong cross-generator generalization, with mIoU consistently above 0.9 and 0.8, respectively. Notably, both models were not specifically designed for such generalization, suggesting that segmentation-based methods hold great potential for transferable AIGC detection.
@@
Segmentation models as strong generalizable whole-image classifiers. Given the strong performance of SegFormer and DeepLabv3+ in binary segmentation, we further investigate whether they can serve as powerful whole-image AIGC classifiers. Unlike standard AIGC classification, which determines whether an entire image is generated by a diffusion model, our task is more challenging: models must identify whether diffusion-edited regions exist within an image. We set a threshold such that if the non-background mask area exceeds a minimal threshold 1% of the image area, the image is considered to contain edited content. We compare with two established AIGC classification models, CNNSpot [cnnspot] and UniversalFakeDet [unifakedet]. During training, we additionally collect 30k real images from COCO [coco] as negative samples to train the classification baselines. Results in Table 5 reveal a striking observation: when repurposed for whole-image classification, segmentation models outperform dedicated AIGC classifiers by a significant margin. We attribute this advantage to the fine-grained, per-pixel supervision employed during segmentation training. Our dataset includes images containing multiple, interleaved edited regions from different generators, allowing segmentation models to develop a deeper understanding of localized generative cues. Notably, although UniversalFakeDet was originally designed for cross-model generalization in whole-image classification, it performs markedly worse than general-purpose segmentation models in cross-model evaluation (Table 6). Cross-model generalization has long been a challenge in the AIGC classification community. Through this experiment, we demonstrate that segmentation models trained on multi-turn edited data exhibit strong potential to address this issue, suggesting a paradigm shift from AIGC classification toward more robust AIGC segmentation.
@@ 6 Relation to global editing
A line of work focuses on global image editing [tan2025ominicontrol, tan2025ominicontrol2, nanobanana2025, wei2024omniedit], where the entire output image is synthesized from scratch. From the perspective of AIGC detection, these methods can be grouped together with standard (“vanilla”) image generation approaches, as they lack the notion of a local editing region—all pixels are fully regenerated.
@@ 7 Do LoRAs affect localization?
Diffusion models can be enhanced with Low-Rank Adaptation (LoRA) modules [hu2022lora] to improve efficiency [ren2024hypersd] or expressiveness [zhang2025easycontrol, song2024processpainter]. In this experiment, we investigate whether the use of LoRAs impacts the localization performance of baseline segmentation models. Specifically, we apply the Hyper-SD LoRA [ren2024hypersd] (8 steps, DDIM scheduler [song2020denoising]) to SDXL [podell2023sdxl], which reduces the number of inference steps from 50 to 8. We re-edit all regions in the validation set that were previously modified by SDXL using the LoRA-accelerated version (SDXL-LoRA). Then, we re-evaluate the baseline semantic segmentation model (trained on 8 vanilla generators from the main text, which just saw SDXL-edited images during training). In this setup, edits made by SDXL-LoRA are considered correctly predicted if classified as SDXL, allowing us to evaluate the segmentation model’s generalization ability to LoRA-based variants. As shown in Table 7, the segmentation model exhibits strong generalization to LoRA variants, with only a minor mIoU drop of approximately 0.03.
@@ 8 Quality assessment
To remove severely low-quality outputs generated by the automated editing framework, we employ Qwen2.5-VL [Qwen2.5-VL] to evaluate the quality of each edited image. Figure 7 shows two chain-of-thought demonstration examples provided to Qwen2.5-VL. Images receiving a score below 3 are discarded, resulting in the removal of roughly 50% of the generated samples. This threshold is intentionally strict, leading to a high rejection rate, but it ensures a higher-quality resulting dataset.
Figure 8 shows typical low-quality edits, usually caused by insufficient diffusion editing capability or inaccurate masks. In the third example, we observe that Qwen2.5-VL occasionally assigns quality scores based on the entire image rather than the edited region, which may inadvertently filter out some higher-quality edits. However, this effect is minor in our dataset construction, as the majority of low-quality edits are successfully removed.
@@ 9 Segmentation on different base images
Here we provide more detailed results on the real- and AI-based subsets. As shown in Tab. 8 Deeplabv3+ performs generally better on AI-based images than on real-image bases.

### v2

@@ 1 Introduction
BR-Gen [brgen] General 1 3 SDXL [podell2023sdxl], BrushNet [ju2024brushnet], PowerPaint [powerpaint]
@@
In this paper, we introduce DiffSeg30k, a publicly available dataset comprising 30k diffusion-edited images with pixel-level annotations, specifically aimed at facilitating the detection and localization of diffusion-based edits. To reflect realistic, practical scenarios, DiffSeg30k incorporates four key features: 1) In-the-wild images—we collect natural images from COCO dataset [coco] to ensure real-world content diversity. In addition, we leverage COCO-derived prompts with diffusion models to generate complementary AI-based images; 2) Diverse diffusion models—local editing is conducted using eight state-of-the-art diffusion models; 3) Realistic user-driven edits—an automated annotation pipeline leveraging vision-language models (VLMs) [Qwen2.5-VL] identifies semantically meaningful regions and generates context-aware editing prompts, including object addition, removal, and attribute modification; and 4) Multi-turn editing—each image undergoes up to three sequential edits using different diffusion models, simulating realistic scenarios in which online-shared images are iteratively modified (see LABEL:fig:teaser). Beyond pixel-level annotations of edited areas, the specific diffusion model used for each edit is annotated, defining a novel semantic segmentation task for fine-grained localization that simultaneously identifies the edited regions and the corresponding editing models. Tab. 1 compares DiffSeg30k with existing benchmarks containing local editing data.
@@ 2.2 AIGC Detection
The task of AIGC detection involves identifying content created by generative AI models such as GANs and diffusion models. Current datasets [zhu2023genimage, defake, bird2024cifake, verdoliva20222022, deepart, he2021forgerynet, guo2023hierarchical, miao2025ddl] and methods [cozzolino2024raising, cnnspot, zhong2023patchcraft, cozzolino2024zero, wang2025scaling, jiang2025edittrack, sun2025fragfake] predominantly focus on classifying entire images. Existing datasets can be categorized into two types according to their image content: facial images [he2021forgerynet, yang2019exposing, wang2019fakespotter, dang2020detection, gandhi2020adversarial, miao2025ddl] and general images [zhu2023genimage, defake, bird2024cifake, verdoliva20222022, deepart]. Several datasets [guillaro2023trufor, wfake, epstein2023online, liu2024hfnet] include limited numbers ($\leq 3$) of diffusion models and only involving single local edit per image, as detailed in Tab. 1. In contrast, our proposed DiffSeg30k dataset focuses on general images, covers eight different diffusion models, and incorporates multi-turn editing by different models on each image. This comprehensive dataset provides systematic support for more general semantic segmentation tasks beyond conventional first-classify-then-binary-segmentation paradigm [guo2023hierarchical, yan2025gamma] used in single-turn editing datasets.
@@ 3.1 Dataset construction
We developed an automated image editing and annotation pipeline based on vision-language models (VLM), illustrated in Fig. 2. The pipeline comprises two stages: 1) generating candidate editing regions and 2) creating suitable editing prompts for these regions, followed by image editing.
@@
Stage 1—generating candidate editing regions. As shown in Fig. 2, given an arbitrary image, we first use Qwen2.5-VL [Qwen2.5-VL] to identify distinct objects present in the image. The identified object categories are then provided to Grounded-SAM [groundedsam] to generate corresponding masks for each object. Subsequently, we calculate the Intersection-over-Union (IoU) among object masks and eliminate redundant masks with IoU greater than 70% to avoid repetitive editing of the same object. Finally, we randomly select between one to three object masks for editing.
@@
Stage 2—editing prompt generation & inpainting. As depicted in Fig. 2, each selected object mask from Stage 1 is paired with the original image and fed into the Qwen2.5-VL model, prompting it to generate contextually appropriate editing prompts. For inpainting, the generated editing prompt, corresponding object mask, and the original image are simultaneously provided to a diffusion model to perform localized inpainting. For multi-turn editing, the edited image can form a new pair with the remaining object masks, repeating the prompt generation and inpainting processes. Thus, we obtain images edited sequentially up to three times, depending on the number of masks selected in Stage 1.
@@
Data diversity & balance. To ensure the diversity and balance of the edited images, we specifically considered the following six aspects: 1) Balanced Candidate Object Types: We prompted the VLM to prioritize selecting humans, as human-centric edits are significant in practical applications, yet VLMs tend to under-select humans naturally. 2) Balanced Candidate Mask Area: We encouraged VLM to prioritize larger objects for editing, counteracting the empirical tendency of VLM to select smaller objects. 3) Balanced Edit Types: During editing prompt generation, we prompted VLM to randomly choose from three edit types: attribute changes, object additions, and object removals. For object additions, we followed the approach of SEED-Data-Edit [ge2024seed], first removing an object and then adding a new object in the same position to ensure natural placement. 4) Balanced Edit Models: For inpainting, we randomly selected one diffusion model from eight state-of-the-art diffusion models to execute an editing prompt. 5) Enhanced Multi-turn Editing: We manually increased the frequency of multi-turn editing, with the first, second, and third turns occurring at an approximate 1:4:5 ratio, as multi-turn editing subsumes single-turn cases and presents a more challenging scenario. 6) Balanced Real and AI Bases: About half of the base images were real, collected from COCO, while the other half were AI-generated using COCO prompts with random diffusion models.
@@ 3.3 Dataset statistics
We conducted a statistical analysis of DiffSeg30k, illustrated in Fig. 3. It reveals that smaller objects are edited more frequently, whereas larger objects tend to be edited less often, with editing frequency smoothly decreasing as object area increases. Objects such as food, vehicles, furniture, nature scenery, animals, humans, and signs are among the most frequently edited. The dataset maintains a balanced distribution across editing models and base image distribution. Fig. 4 shows some editing results.
@@ 4.1 Experimental setup
Training details. Due to the lack of existing baseline segmentation models on general AIGC localization to handle semantic segmentation, we trained two CNN baselines based on classic semantic segmentation frameworks: FCN-8s [long2015fully] and Deeplabv3+[chen2017rethinking, chen2018encoder] with ResNet50[he2016deep] and 1 transfomer baseline SegFormer-B2 [xie2021segformer]. We evaluated these models on two tasks: binary segmentation (localizing edited areas) and semantic segmentation (localizing edited areas while distinguishing among different editing models). The dataset is split into training and validation sets in an 8:2 ratio, and all results are reported on the validation set. Models are trained on 512×512 resolution (for acceleration) using default hyperparameters from open-source implementations. We additionally include 5k unedited real COCO images in the training set to help suppress false positives on real images. Training FCN-8s, SegFormer-B2, and Deeplabv3+ takes approximately 8, 6, and 3 hours, respectively, on an NVIDIA 4090 GPU.
@@ 4.2 Results
Model capacity matters. As shown in Tab. 2, the older FCN-8s struggles with both binary and semantic localization of diffusion-edited regions. In contrast, more powerful segmentation models Deeplabv3+ and SegFormer , performs substantially better, achieving strong binary localization with an mIoU of 0.974 and 0.961. These results highlight the importance of model representational capacity.
Semantic segmentation is challenging. Simultaneously localizing edited regions and attributing them to the correct diffusion model is substantially more difficult than binary localization. Even for the best performed model SegFormer, the mIoU for semantic segmentation drops to 0.825, leaving significant room for improvement. For Deeplabv3+ and FCN, this number further drops significantly to 0.76 and 0.23. Fig. 5 shows the confusion matrices: FCN-8s tends to misclassify edits from all models as background pixels, while SegFormer struggles particularly with Flux. Deeplab v3+ struggles with Kolors and SDXL. It also often misclassifies SD2 and Flux. Fig. 6 further illustrates representative results. FCN-8s produces largely random outputs in both tasks, whereas SegFormer and Deeplab v3+ achieves more coherent segmentations but occasionally misclassifies the editing model. In the following detailed ablation, we only evaluate Deeplabv3+ and SegFormer.
Baseline models are sensitive to image transformations. We evaluated the robustness of SegFormer and DeepLabv3+ under common image transformations, including JPEG compression and resizing. As shown in Tab. 3, image transformations substantially degraded the performance of both Deeplabv3+ and SegFormer, with severe mIoU drops—semantic segmentation almost failed under JPEG compression. This highlights the need for improved network architectures or data augmentation strategies to achieve deployable localization models. Interestingly, DeepLabv3+ exhibits partial resilience in binary segmentation settings, while SegFormer demonstrates certain robustness when resized to higher resolutions.
Segmentation models achieve strong cross-generator generalization. Tab. 4 presents the cross-generator generalization results of SegFormer and Deeplabv3+. Training was conducted on images edited by six diffusion models and tested on edits from two unseen models, with three different combinations evaluated. Both SegFormer and Deeplabv3+ demonstrate strong cross-generator generalization, with mIoU consistently above 0.9 and 0.8, respectively. Notably, both models were not specifically designed for such generalization, suggesting that segmentation-based methods hold great potential for transferable AIGC detection.
@@
Segmentation models as strong generalizable whole-image classifiers. Given the strong performance of SegFormer and DeepLabv3+ in binary segmentation, we further investigate whether they can serve as powerful whole-image AIGC classifiers. Unlike standard AIGC classification, which determines whether an entire image is generated by a diffusion model, our task is more challenging: models must identify whether diffusion-edited regions exist within an image. We set a threshold such that if the non-background mask area exceeds a minimal threshold 1% of the image area, the image is considered to contain edited content. We compare with two established AIGC classification models, CNNSpot [cnnspot] and UniversalFakeDet [unifakedet]. During training, we additionally collect 30k real images from COCO [coco] as negative samples to train the classification baselines. Results in Tab. 5 reveal a striking observation: when repurposed for whole-image classification, segmentation models outperform dedicated AIGC classifiers by a significant margin. We attribute this advantage to the fine-grained, per-pixel supervision employed during segmentation training. Our dataset includes images containing multiple, interleaved edited regions from different generators, allowing segmentation models to develop a deeper understanding of localized generative cues. Notably, although UniversalFakeDet was originally designed for cross-model generalization in whole-image classification, it performs markedly worse than general-purpose segmentation models in cross-model evaluation (Tab. 6). Cross-model generalization has long been a challenge in the AIGC classification community. Through this experiment, we demonstrate that segmentation models trained on multi-turn edited data exhibit strong potential to address this issue, suggesting a paradigm shift from AIGC classification toward more robust AIGC segmentation.
@@ 6 Relation to global editing
A line of work focuses on global image editing [tan2025ominicontrol, tan2025ominicontrol2, nanobanana2025, wei2024omniedit, cocoinpaint], where the entire output image is synthesized from scratch. From the perspective of AIGC detection, these methods can be grouped together with standard (“vanilla”) image generation approaches, as they lack the notion of a local editing region—all pixels are fully regenerated.
@@ 7 Do LoRAs affect localization?
Diffusion models can be enhanced with Low-Rank Adaptation (LoRA) modules [hu2022lora] to improve efficiency [ren2024hypersd] or expressiveness [zhang2025easycontrol, song2024processpainter]. In this experiment, we investigate whether the use of LoRAs impacts the localization performance of baseline segmentation models. Specifically, we apply the Hyper-SD LoRA [ren2024hypersd] (8 steps, DDIM scheduler [song2020denoising]) to SDXL [podell2023sdxl], which reduces the number of inference steps from 50 to 8. We re-edit all regions in the validation set that were previously modified by SDXL using the LoRA-accelerated version (SDXL-LoRA). Then, we re-evaluate the baseline semantic segmentation model (trained on 8 vanilla generators from the main text, which just saw SDXL-edited images during training). In this setup, edits made by SDXL-LoRA are considered correctly predicted if classified as SDXL, allowing us to evaluate the segmentation model’s generalization ability to LoRA-based variants. As shown in Tab. 7, the segmentation model exhibits strong generalization to LoRA variants, with only a minor mIoU drop of approximately 0.03.
@@ 8 Quality assessment
To remove severely low-quality outputs generated by the automated editing framework, we employ Qwen2.5-VL [Qwen2.5-VL] to evaluate the quality of each edited image. Fig. 7 shows two chain-of-thought demonstration examples provided to Qwen2.5-VL. Images receiving a score below 3 are discarded, resulting in the removal of roughly 50% of the generated samples. This threshold is intentionally strict, leading to a high rejection rate, but it ensures a higher-quality resulting dataset.
Fig. 8 shows typical low-quality edits, usually caused by insufficient diffusion editing capability or inaccurate masks. In the third example, we observe that Qwen2.5-VL occasionally assigns quality scores based on the entire image rather than the edited region, which may filter out some higher-quality edits. However, this does not harm the overall dataset quality, as our guiding principle is to err on the side of stricter filtering—prefer removing a few good samples rather than keeping any low-quality ones.
@@ 9 Segmentation on different base images
Here we provide more detailed results on the real- and AI-based subsets. As shown in Tab. 8, Deeplabv3+ performs generally better on AI-based images than on real-image bases.